OpenStudy (anonymous):

Graph this function and identify the local minimum and maximum: f(x)=2x^2-x^4

OpenStudy (anonymous):

get the first derivative equate to 0

OpenStudy (campbell_st):

you will need to find the 1st and 2nd derivatives \[f(x) = 2x^2 - x^4\] \[f'(x) = 4x - 4x^3\] \[f"(x) = 4 - 12x^2\] solve f'(x) = 0 this will give the stationary points \[4x - 4x^3 = 0\] \[4x(1 -x)(1+x) = 0\] x = -1, 0, 1 these the stationary points in the 2nd derivative to determine maximum, minimum or points of inflexion f"(x) > 0 minimum f"(x) < 0 maximum f"(0) = 0 point of inflexion

OpenStudy (anonymous):

"these the stationary points in the 2nd derivative to determine maximum, minimum or points of inflexion f"(x) > 0 minimum f"(x) < 0 maximum f"(0) = 0 point of inflexion" I don't understand this, can you explain it again?

OpenStudy (campbell_st):

well the stationary points from the 1st derivative are x = -1, 0, 1 the find f"(-1) = 4 - 12(-1)^2 = - 8 then a maximum exists at (-1, 1) since f"(-1) < 0 next f"(0) = 4 - 12(0)^2 = 4 therefore a minumum at (0, 0) since f"(0) > 0 and f"(1) = 4 - 12(1)^2 = -8 therefore a maximum since f"(x) > 0 its the best way to check stationary points.... the alternative method is to pick points close to the stationary points and check for a change in the gradients -ve, 0 , +ve is a minimum and +ve, 0, -ve is a maximum  the points of inflexion occur ( chnage in concavity) where f"(x) = 0 or 0 = 4 - 12x^2 \[ x = \pm \frac{1}{\sqrt{3}}\]